# RaCalNet: Radar Calibration Network for Sparse-Supervised Metric Depth Estimation

### v1

@@ Abstract
Dense metric depth estimation using millimeter-wave radar typically requires dense LiDAR supervision, generated via multi-frame projection and interpolation, to guide the learning of accurate depth from sparse radar measurements and RGB images. However, this paradigm is both costly and data-intensive. To address this, we propose RaCalNet, a novel framework that eliminates the need for dense supervision by using sparse LiDAR to supervise the learning of refined radar measurements, resulting in a supervision density of merely around 1% compared to dense-supervised methods. Unlike previous approaches that associate radar points with broad image regions and rely heavily on dense labels, RaCalNet first recalibrates and refines sparse radar points to construct accurate depth priors. These priors then serve as reliable anchors to guide monocular depth prediction, enabling metric-scale estimation without resorting to dense supervision. This design improves structural consistency and preserves fine details. Despite relying solely on sparse supervision, RaCalNet surpasses state-of-the-art dense-supervised methods, producing depth maps with clear object contours and fine-grained textures. Extensive experiments on the ZJU-4DRadarCam dataset and real-world deployment scenarios demonstrate its effectiveness, reducing RMSE by 35.30% and 34.89%, respectively.
@@ I Introduction
Millimeter-wave (mmWave) radar provides direct metric measurements and robust performance under diverse lighting and texture variations, making it a viable sensing modality for depth completion in autonomous systems. However, mmWave point clouds are extremely sparse, often two to three orders of magnitude sparser than LiDAR, and contaminated by noise, which makes the generation of dense, high-fidelity depth maps highly challenging [1].
@@
Current radar-guided depth estimation methods typically rely on LiDAR supervision to guide depth completion. They associate radar points with image pixels to generate semi-dense priors, which are then processed by completion networks to produce dense depth maps. Crucially, these methods rely on dense LiDAR supervision synthesized via multi-frame accumulation and interpolation. Singh [3] introduced this paradigm by aggregating 161 consecutive frames (80 past, 80 future, and the current frame) to construct dense supervision signals, a practice subsequently adopted in [2]. However, this strategy suffers from three critical limitations: (i) high data collection and storage costs due to the need for extensive LiDAR scans; (ii) interpolation artifacts that degrade boundary sharpness and structural accuracy; and (iii) susceptibility to dynamic scene misalignments, which introduce inconsistency and noise in the supervision signal.
@@
To address these limitations, we propose RaCalNet, a novel radar-guided depth estimation framework that eliminates the need for dense LiDAR supervision by refining sparse radar points into reliable geometric anchors. Specifically, we first train a cross-modal attention module using only single-frame sparse LiDAR to learn feature correlations between radar and RGB inputs. This module predicts confidence weights for radar points, effectively suppressing noise and refining their pixel-level projections through a lightweight displacement prediction head. The resulting refined radar points are then employed as geometric anchors to calibrate monocular depth predictions via an optimization framework. By minimizing a geometric alignment loss, we solve for global scale and shift parameters that align predicted depths with radar geometry, ensuring metric consistency with physical measurements. As shown in Fig. 1, while Radarcam [2] relies on dense LiDAR supervision, our method achieves superior depth estimation using only sparse LiDAR supervision ($\sim$1point density), yielding sharper structural boundaries and significantly reduced scale errors.
@@
Figure 2: Overview of our proposed RaCalNet framework. The system consists of two main components: (i) a Radar Refinement Module, which performs radar point screening and pixel-wise displacement refinement; and (ii) a Metric Depth Optimizer, which fuses refined radar with monocular depth for metric-accurate output.
@@
Our key contributions are summarized as follows: (i) We present a novel radar-guided depth estimation framework, RaCalNet, which eliminates the need for dense LiDAR supervision, replacing it with a recalibration and optimization strategy guided by refined radar points; (ii) we leverage a new radar refinement module with cross-modal attention to identify reliable radar points and recalibrate their pixel-level projections, generating sparse yet precise reference anchors for downstream depth prediction; (iii) We design a LiDAR-free optimization strategy that aligns monocular predictions with refined radar anchors, producing dense depth maps with accurate scale and sharp structural details; (iv) We validate our approach on the public ZJU-4DRadarCam dataset and real-world deployment scenarios, demonstrating significant performance gains over existing methods, including up to 35.30% RMSE reduction.
@@ II-A Depth Estimation with Monocular Vision
To address this, various strategies introduce auxiliary cues. VISLAM [11] performs geometric rectification of camera perspectives, while CodeVIO [9] jointly optimizes visual-inertial odometry (VIO) and dense depth through coupled estimation. Xie et al. [10] further enhance accuracy via flow-depth consistency and multi-frame triangulation. More recently, Wofk et al. [8] propose a metric depth estimation framework that globally aligns monocular predictions with sparse but accurate VIO measurements, followed by local optimization. While effective, this approach still relies on tightly coupled VIO systems and may suffer from drift or degraded performance in textureless environments. In parallel, a series of video-based methods [12, 13, 14, 15, 16] enhance temporal consistency through learned regularization across frames. These techniques are largely orthogonal to our work and can benefit from our approach as a strong initialization stage.
While these methods significantly improve depth prediction quality, they either rely on tightly coupled VIO systems or require temporal sequences and large-scale supervision, limiting their applicability in unconstrained or dynamic environments. In contrast, our goal is to improve monocular depth accuracy by introducing sparse yet metrically accurate millimeter-wave radar as geometric anchors. Unlike VIO or dense LiDAR, radar offers robust and cost-effective metric signals that are resilient to lighting and texture variations.
@@ II-B Depth Estimation from Radar-Camera Fusion
To address this, prior works have explored various fusion strategies. Lin et al. [17] introduced a two-stage encoder–decoder pipeline to denoise radar point clouds before fusing them with image features. Lo et al. [18] proposed height-extended radar representations to enhance spatial context. Long et al. [1] designed a Radar-to-Pixel (R2P) network that leverages Doppler velocity and optical flow to establish pixel-level correspondences between radar and image domains. Their follow-up work [19] integrated multi-sweep radar accumulation with dense LiDAR supervision in an image-guided completion framework.
@@ III METHOD
In this work, we address the task of dense, metrically accurate depth estimation $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ by fusing monocular RGB images $I\in\mathbb{R}^{3\times H\times W}$ with sparse radar data $P=\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{k}$. The overall architecture of our RaCalNet is illustrated in Fig. 2. Our pipeline first recalibrates and refines radar measurements using sparse LiDAR supervision in Radar Refinement Module, producing accurate depth anchors. These refined anchors are then used to optimize monocular depth predictions, specifically to correct their scale and enhance their structural consistency, resulting in accurate metric-scale estimation. Each component of our framework is detailed in the following subsections.
III-A Radar Screening and Pixel-wise Displacement Refinement
@@
The radar recalibration module employs a cross-modal fusion design. The image encoder is based on ResNet-34 [21], producing hierarchical feature maps with channel dimensions of 32, 64, 128, 256, and 512. These multi-scale features are aggregated using a Feature Pyramid Network (FPN), ensuring the preservation of spatial detail and semantic richness. To further enhance feature alignment and interaction, we incorporate channel and spatial attention mechanisms, which help highlight salient structures and suppress irrelevant background noise in a content-adaptive manner.
The radar encoder is implemented as a multi-layer perceptron (MLP) with fully connected layers of dimensions 32, 64, 128, 256, and 512. To facilitate effective radar-image fusion, image features $\mathbf{F}_{img}\in\mathbb{R}^{B\times K\times D\times H\times W}$ are reshaped along the spatial dimensions to $\mathbb{R}^{B\times K\times D\times N_{img}}$ with $N_{img}=H\times W$, so as to align with radar features $\mathbf{F}_{rad}\in\mathbb{R}^{B\times K\times D}$, where $B$ is the batch size, $K$ is the number of radar points, and $D$ is the feature dimension. Both modalities are then flattened along the spatial axis and passed through stacked attention layers, forming the CamRad-Fusion module for effective cross-modal interaction and feature enhancement.
@@
where $\mathbf{Q}=\mathbf{X}_{rad}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}_{img}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{X}_{img}\mathbf{W}_{V}$. The output preserves the radar feature resolution and is reshaped back to $\mathbb{R}^{B\times K\times D}$ for subsequent processing. This cross-modal attention allows the network to enhance radar representations with informative visual cues. Finally, the refined radar features are fed into the refinement head, which consists of a two-layer MLP that separately predicts a confidence score and a pixel-wise displacement for each radar point.
@@
3) Training: The network is trained using sparse LiDAR points projected onto the image plane as depth ground truth, denoted as $d_{gt}(x,y)$ for valid LiDAR pixel locations. Dual learning objectives are defined over the radar points $p\in P$ for confidence classification and displacement regression.
@@ III-B Metric Depth Optimization
Let $d_{\text{radar}}^{(i)}$ denote the refined metric depth of radar point $p_{i}$, and $(u_{i},v_{i})$ its corresponding image coordinates. To align these measurements with the predicted $\hat{d}_{m}$, we solve the following least squares problem to find optimal scale $\alpha$ and shift $\beta$:
$\min_{\alpha,\beta}\sum_{i\in\mathcal{S}}\left(\frac{1}{d_{radar}^{(i)}}-\alpha\cdot{\hat{d}_{m}(u_{i},v_{i})}-\beta\right)^{2}$ (11)
where $\mathcal{S}$ denotes the subset of radar points used for alignment. To ensure robustness, we discard outliers by filtering invalid or unreliable regions in $\hat{d}_{m}$.
Instead of using a fixed threshold, we adopt a dynamic threshold selection strategy. Specifically, we define a candidate set of inverse depth thresholds $\mathcal{T}_{d}=\{t_{1},t_{2},\dots,t_{M}\}$, and determine the best threshold $T^{*}$ by selecting the one that yields the lowest fitting error. For each threshold $t\in\mathcal{T}_{d}$, we define the corresponding valid set:
$\mathcal{S}_{t}=\left\{i\mid 0<\hat{d}_{m}(u_{i},v_{i})<t\right\}$ (12)
and compute its associated parameters $(\alpha_{t},\beta_{t})$ using Equation (11). The optimal threshold $T^{*}$ and corresponding parameters are selected as:
$T^{*}=\arg\min_{t\in\mathcal{T}_{d}}\sum_{i\in\mathcal{S}_{t}}\left(\frac{1}{d_{\text{radar}}^{(i)}}-\alpha_{t}\cdot\hat{d}_{m}(u_{i},v_{i})-\beta_{t}\right)^{2}$ (13)
Finally, the recovered metric depth map $\hat{d}$ is obtained by inverting the affine-aligned inverse depth:
$\hat{d}(u,v)=\begin{cases}\frac{1}{\alpha^{*}\cdot\hat{d}_{m}(u,v)+\beta^{*}}&\text{if }\hat{d}_{m}(u,v)\neq 0\\ \text{undefined}&\text{otherwise}\end{cases}$ (14)
Note that Equation (14) leaves regions with zero inverse depth undefined, which typically correspond to sky or invalid areas. This design enables flexible handling for downstream applications, where such regions can be explicitly filled using task-specific strategies (e.g., assigning 0 or a large constant). To summarize the process, the optimization pipeline is outlined in Algorithm 1.
Algorithm 1 Global Metric Depth Optimization
0:  Monocular inverse depth map $\hat{d}_{m}$, radar points $\{(u_{i},v_{i},d_{\text{radar}}^{(i)})\}_{i=1}^{N}$, threshold set $\mathcal{T}_{d}$
0:  Optimized depth map $\hat{d}$
1:  Initialize $L^{*}\leftarrow+\infty$, $T^{*}\leftarrow 0$
@@
3:     $\mathcal{S}_{t}\leftarrow\{i\mid 0<\hat{d}_{m}(u_{i},v_{i})<t\}$
4:     Fit $(\alpha_{t},\beta_{t})$ by minimizing:
$\sum_{i\in\mathcal{S}_{t}}\left(\frac{1}{d_{\text{radar}}^{(i)}}-\alpha\cdot\hat{d}_{m}(u_{i},v_{i})-\beta\right)^{2}$
5:     Compute $L_{t}$
6:     if $L_{t}<L^{*}$ then
7:        $L^{*}\leftarrow L_{t}$, $T^{*}\leftarrow t$, $(\alpha^{*},\beta^{*})\leftarrow(\alpha_{t},\beta_{t})$
8:     end if
9:  end for
10:  for all $(u,v)$ such that $\hat{d}_{m}(u,v)\neq 0$ do
11:     $\hat{d}(u,v)\leftarrow\frac{1}{\alpha^{*}\cdot\hat{d}_{m}(u,v)+\beta^{*}}$
12:  end for
13:  return  $\hat{d}$

### v2

# RaCalNet: Radar Calibration Network for Sparse-Supervised Dense Depth Estimation

@@ Abstract
Dense depth estimation using millimeter-wave radar typically requires dense LiDAR supervision, generated via multi-frame projection and interpolation, for guiding the learning of accurate depth from sparse radar measurements and RGB images. However, this paradigm is both costly and data-intensive. To address this, we propose RaCalNet, a novel framework that eliminates the need for dense supervision by using sparse LiDAR to supervise the learning of refined radar measurements, resulting in a supervision density of merely around 1% compared to dense-supervised methods. RaCalNet is composed of two key modules. The Radar Recalibration module performs radar point screening and pixel-wise displacement refinement, producing accurate and reliable depth priors from sparse radar inputs. These priors are then used by the Metric Depth Optimization module, which learns to infer scene-level scale priors and fuses them with monocular depth predictions to achieve metrically accurate outputs. This modular design enhances structural consistency and preserves fine-grained geometric details. Despite relying solely on sparse supervision, RaCalNet produces depth maps with clear object contours and fine-grained textures, demonstrating superior visual quality compared to state-of-the-art dense-supervised methods. Quantitatively, it achieves performance comparable to existing methods on the ZJU-4DRadarCam dataset and yields a 34.89% RMSE reduction in real-world deployment scenarios. We plan to gradually release the code and models in the future at https://github.com/818slam/RaCalNet.git.
Index Terms:
Depth estimation, Millimeter-wave radar, Sensor fusion, Sparse supervision, Monocular vision
@@ I Introduction
Millimeter-wave (mmWave) radar has been increasingly adopted in modern measurement systems for its ability to provide direct metric observations under diverse environmental conditions. In particular, radar has enabled effective solutions in diverse tasks such as displacement estimation [1, 2], simultaneous localization and mapping (SLAM) [3, 4], object detection [5, 6], and gesture recognition [7, 8]. These capabilities highlight its growing relevance in instrumentation scenarios where accurate and robust perception of physical phenomena is required.
@@
Unlike conventional 3D mmWave radar, which only provides azimuth and range measurements, 4D mmWave radar introduces elevation sensing, resulting in denser point clouds with full spatial information. This richer representation has positioned 4D radar as a promising alternative to LiDAR for various perception tasks. However, mmWave radar point clouds-despite the added vertical dimension-remain extremely sparse and noisy, often two to three orders of magnitude sparser than LiDAR [10], posing critical challenges for downstream tasks such as dense depth estimation, which is fundamental for 3D reconstruction and autonomous navigation.
To overcome radar sparsity, existing methods often associate radar points with image pixels to create sparse depth priors, which are then densified using learning-based completion networks [9, 11]. To supervise training, they typically synthesize dense labels by aggregating multi-frame LiDAR scans and interpolating missing regions. However, such practices not only incur heavy data collection and processing overhead but also introduce spatial artifacts that degrade depth accuracy and geometric fidelity.
@@
To address these limitations, we propose RaCalNet, a novel radar-guided depth estimation framework that eliminates the need for dense LiDAR supervision by refining sparse radar points into reliable geometric anchors. Specifically, the Radar Recalibration module leverages a cross-modal attention mechanism, trained with single-frame sparse LiDAR, to learn feature correlations between radar and RGB inputs. This module predicts confidence weights for radar points, effectively suppressing noise and refining their pixel-level projections through a lightweight displacement prediction head. The resulting refined radar points serve as geometric anchors for the Metric Depth Optimization module, which calibrates monocular depth predictions through a multi-stage process involving global alignment, cluster-wise refinement, and edge-aware smoothing. This module minimizes a geometric alignment loss to enforce consistency between predicted depths and radar geometry, enabling metrically accurate depth estimation aligned with physical measurements. As shown in Fig. 1, while Radarcam [9] relies on dense LiDAR supervision, our method achieves superior depth estimation using only sparse LiDAR supervision ($\sim$1point density), yielding sharper structural boundaries and significantly reduced scale errors.
@@
Figure 2: Overview of our proposed RaCalNet framework. The system consists of two main components: (i) a Radar Recalibration module, which performs radar point screening and pixel-wise displacement refinement; and (ii) a Metric Depth Optimizer, which fuses refined radar with monocular depth for metric-accurate output.
@@
Our key contributions are summarized as follows: (i) We present a novel radar-guided depth estimation framework, RaCalNet, which eliminates the need for dense LiDAR supervision, replacing it with a recalibration and optimization strategy guided by refined radar points; (ii) We leverage a new Radar Recalibration module with cross-modal attention to identify reliable radar points and recalibrate their pixel-level projections, generating sparse yet precise reference anchors for downstream depth prediction; (iii) We design a Metric Depth Optimization module that aligns monocular predictions with refined radar anchors, producing dense depth maps with accurate scale and sharp structural details; (iv) We validate our approach on the public ZJU-4DRadarCam dataset and real-world deployment scenarios, demonstrating significant performance gains over existing methods, including up to 35.30% RMSE reduction.
@@ II-A Depth Estimation with Monocular Vision
To address this, various strategies introduce auxiliary cues. VISLAM [19] performs geometric rectification of camera perspectives, while CodeVIO [17] jointly optimizes visual-inertial odometry (VIO) and dense depth through coupled estimation. Yin et al. [20] proposed an unsupervised monocular depth estimation method that fuses inertial and visual features to produce scale-consistent depth for SLAM systems. More recently, Wofk et al. [16] propose a metric depth estimation framework that globally aligns monocular predictions with sparse but accurate VIO measurements, followed by local optimization. While effective, this approach still relies on tightly coupled VIO systems and may suffer from drift or degraded performance in textureless environments. Yuan et al. [21] proposed a self-supervised monocular depth framework with a depth-motion prior to guide pseudo-LiDAR generation without external calibration. In parallel, a series of video-based methods [22, 23, 24, 25, 26] enhance temporal consistency through learned regularization across frames.
While these methods significantly improve depth prediction quality, they either rely on tightly coupled VIO systems or require temporal sequences and large-scale supervision, limiting their applicability in unconstrained or dynamic environments. In contrast, our goal is to improve monocular depth accuracy by introducing sparse yet metrically accurate millimeter-wave radar as geometric anchors.
@@ II-B Depth Estimation from Radar-Camera Fusion
To address this, prior works have explored various fusion strategies. Lin et al. [27] introduced a two-stage encoder–decoder pipeline to denoise radar point clouds before fusing them with image features. Lo et al. [28] proposed height-extended radar representations to enhance spatial context. Long et al. [10] designed a Radar-to-Pixel (R2P) network that leverages Doppler velocity and optical flow to establish pixel-level correspondences between radar and image domains. Their follow-up work [29] integrated multi-sweep radar accumulation with dense LiDAR supervision in an image-guided completion framework. Cui et al. [30] fused event camera data with sparse depth using density-based point clustering and surface reconstruction to generate dense depth maps.
@@ III Method
In this work, we address the task of dense, metrically accurate depth estimation $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ by fusing monocular RGB images $I\in\mathbb{R}^{3\times H\times W}$ with sparse radar data $P=\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{k}$. The overall architecture of our RaCalNet is illustrated in Fig. 2. Our pipeline first recalibrates and refines radar measurements using sparse LiDAR supervision in Radar Recalibration module, producing accurate depth anchors. These refined anchors are then used to optimize monocular depth predictions, specifically to correct their scale and enhance their structural consistency, resulting in accurate metric-scale estimation. Each component of our framework is detailed in the following subsections.
III-A Radar Recalibration Module
@@
Motivated by recent advances in sparse 3D representation and cross-modal point processing [32, 33], our Radar Recalibration module employs a cross-modal fusion design. The image encoder is based on ResNet-34 [34], producing hierarchical feature maps with channel dimensions of 32, 64, 128, 256, and 512. These multi-scale features are aggregated using a Feature Pyramid Network (FPN), ensuring the preservation of spatial detail and semantic richness. To further enhance feature alignment and interaction, we incorporate channel and spatial attention mechanisms, which help highlight salient structures and suppress irrelevant background noise in a content-adaptive manner.
The radar encoder is implemented as a multi-layer perceptron (MLP) with fully connected layers of dimensions 32, 64, 128, 256, and 512. To facilitate effective radar-image fusion, image features $\mathbf{F}_{img}\in\mathbb{R}^{B\times K\times D\times H\times W}$ are reshaped along the spatial dimensions to $\mathbb{R}^{B\times K\times D\times N_{img}}$ with $N_{img}=H\times W$, so as to align with radar features $\mathbf{F}_{rad}\in\mathbb{R}^{B\times K\times D}$, where $B$ is the batch size, $K$ is the number of radar points, and $D$ is the feature dimension. Both modalities are then flattened along the spatial axis and passed through stacked attention layers, forming the CamRad-Fusion net for effective cross-modal interaction and feature enhancement.
@@
where $\mathbf{Q}=\mathbf{X}_{rad}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}_{img}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{X}_{img}\mathbf{W}_{V}$. The output preserves the radar feature resolution and is reshaped back to $\mathbb{R}^{B\times K\times D}$ for subsequent processing. This cross-modal attention allows the network to enhance radar representations with informative visual cues. Finally, the refined radar features are fed into the refinement heads, which consist of a two-layer MLP that separately predicts a confidence score and a pixel-wise displacement for each radar point.
@@
3) Training Loss: The network is trained using sparse LiDAR points projected onto the image plane as depth ground truth, denoted as $d_{gt}(x,y)$ for valid LiDAR pixel locations. Dual learning objectives are defined over the radar points $p\in P$ for confidence classification and displacement regression.
@@ III-B Metric Depth Optimization
Algorithm 1 Metric Depth Optimization
0:  Monocular inverse depth map $\hat{d}_{m}$, radar point set $\{(u_{i},v_{i},d_{\text{radar}}^{(i)})\}$, threshold candidates $\mathcal{T}_{d}$, regularization weight $\lambda$
0:  Refined metric depth $\hat{d}_{\text{final}}$
1:  Stage 1: Global Affine Alignment
@@
3:     $\mathcal{S}_{t}\leftarrow\{i\mid 0<\hat{d}_{m}(u_{i},v_{i})<t\}$Filter unreliable predictions}
4:     $X\leftarrow[\hat{d}_{m}(u_{i},v_{i}),\mathbf{1}]_{i\in\mathcal{S}_{t}}$, $Y\leftarrow[1/d_{\text{radar}}^{(i)}]_{i\in\mathcal{S}_{t}}$
5:     Solve $\min_{\theta}\|Y-X\theta\|^{2}+\lambda\|\theta\|^{2}$, obtain $\theta_{t}=(\alpha_{t},\beta_{t})$
6:     Compute relative error: $\mathcal{L}_{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}\frac{|d_{\text{pred}}^{(i)}-d_{\text{radar}}^{(i)}|}{d_{\text{radar}}^{(i)}}$
7:     if $\mathcal{L}_{t}<\mathcal{L}^{*}$ then
8:        $T^{*}\leftarrow t$, $\theta^{*}\leftarrow\theta_{t}$
9:     end if
10:  end for
11:  Apply global alignment: $\hat{d}_{\text{global}}(u,v)\leftarrow 1/(\alpha^{*}\cdot\hat{d}_{m}(u,v)+\beta^{*})$
12:  Stage 2: Cluster-wise Local Refinement
13:  Extract clustering features:     $\mathbf{f}(u,v)=[u/W,v/H,\log(\hat{d}_{\text{global}}(u,v))]$
14:  Determine number of clusters:     $k\leftarrow\min\left(8,\max\left(2,\left\lfloor N_{\text{valid}}/500\right\rfloor\right)\right)$
15:  Cluster valid pixels via K-means: $\mathcal{C}\leftarrow\{\mathcal{C}_{1},...,\mathcal{C}_{k}\}$
16:  for all cluster $\mathcal{C}_{c}$ do
17:     Identify radar samples within cluster: $\mathcal{S}_{c}\leftarrow\{i\mid(u_{i},v_{i})\in\mathcal{C}_{c}\}$
18:     if $|\mathcal{S}_{c}|\geq 3$ then
19:        $X_{c}\leftarrow[\hat{d}_{m}(u_{i},v_{i}),\mathbf{1}]_{i\in\mathcal{S}_{c}}$, $Y_{c}\leftarrow[1/d_{\text{radar}}^{(i)}]_{i\in\mathcal{S}_{c}}$
20:        Solve $\min_{\theta_{c}}\|Y_{c}-X_{c}\theta_{c}\|^{2}$, obtain $(\alpha_{c},\beta_{c})$
21:        Update: $\hat{d}_{\text{local}}^{(c)}(u,v)\leftarrow 1/(\alpha_{c}\cdot\hat{d}_{m}(u,v)+\beta_{c}),\forall(u,v)\in\mathcal{C}_{c}$
22:     else
23:        Use global prediction: $\hat{d}_{\text{local}}^{(c)}\leftarrow\hat{d}_{\text{global}}$
24:     end if
25:  end for
26:  Stage 3: Edge-Aware Smoothing
27:  Compute gradient magnitude: $E(u,v)\leftarrow\|\nabla\hat{d}_{\text{global}}(u,v)\|$
28:  Define edge mask: $\mathcal{M}_{\text{edge}}\leftarrow\{(u,v)\mid E(u,v)>\tau\}$
29:  Apply Gaussian smoothing: $\hat{d}_{\text{smooth}}\leftarrow\text{GaussianBlur}(\hat{d}_{\text{local}},\sigma=1.5,\text{kernel}=5\times 5)$
30:  Fuse results:
$\hat{d}_{\text{final}}(u,v)\leftarrow\begin{cases}\hat{d}_{\text{local}}(u,v),&(u,v)\in\mathcal{M}_{\text{edge}}\\ \hat{d}_{\text{smooth}}(u,v),&\text{otherwise}\end{cases}$
31:  return  $\hat{d}_{\text{final}}$
To achieve robust alignment with sparse radar measurements, we propose a multi-stage optimization strategy that progressively refines the scale and bias in both global and local contexts. Let $d_{\text{radar}}^{(i)}$ denote the depth of a refined radar point $p_{i}$, and $(u_{i},v_{i})$ its image coordinates. Our approach includes three key components:
1) Global Affine Alignment: We begin by estimating a global affine transformation in the inverse depth space. Given predicted inverse depths $\hat{d}_{m}(u_{i},v_{i})$ and ground-truth inverse depths $1/d_{\text{radar}}^{(i)}$, we solve the following regularized least squares problem:
$\min_{\alpha,\beta}\sum_{i\in\mathcal{S}}\left(\frac{1}{d_{\text{radar}}^{(i)}}-\alpha\cdot\hat{d}_{m}(u_{i},v_{i})-\beta\right)^{2}+\lambda(\alpha^{2}+\beta^{2})$ (11)
where $\mathcal{S}$ denotes the set of valid radar-projected pixels. To enhance robustness, we dynamically search for the optimal inverse depth threshold $T^{*}$ to filter out unreliable predictions:
$\mathcal{S}_{t}=\left\{i\mid 0<\hat{d}_{m}(u_{i},v_{i})<t\right\},\quad T^{*}=\arg\min_{t}\mathcal{L}(t)$ (12)
where $\mathcal{L}(t)$ is the relative alignment error under threshold $t$.
2) Depth-Aware Local Refinement: After global alignment, we further refine the depth map by leveraging local geometric consistency. We perform unsupervised K-means clustering on valid pixels using normalized spatial coordinates and logarithmic aligned depth:
$\mathbf{f}(u,v)=\left[\frac{u}{H},\frac{v}{W},\log(\hat{d}_{\text{global}}(u,v))\right]$ (13)
Each cluster is assumed to represent a locally planar or smoothly varying region. Within each cluster, we identify radar points and estimate a local affine transformation $(\alpha_{c},\beta_{c})$ via least squares, then update depth values:
$\hat{d}_{c}(u,v)=\frac{1}{\alpha_{c}\cdot\hat{d}_{m}(u,v)+\beta_{c}},\quad\forall(u,v)\in\text{cluster }c$ (14)
3) Edge-Aware Smoothing: To enforce spatial smoothness while preserving depth discontinuities, we compute edge maps using Sobel gradients of the aligned depth. A Gaussian filter is then selectively applied to non-edge regions, suppressing high-frequency noise in homogeneous areas without blurring object boundaries.
Algorithm 1 summarizes the full optimization process. This multi-stage alignment improves both the metric accuracy and geometric coherence of monocular depth predictions.